Circular No. 5650

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRO J0422+32 T. Mazeh, E. M. Leibowitz, S. Hemar, and H. Mendelson, Wise Obser-

vatory, Tel Aviv University; and R. E. McCrosky, Oak Ridge Observatory, Harvard-Smithsonian Center for Astrophysics, report: “The optical counterpart of GRO J0422+32 has been monitored at Wise and Oak Ridge Observatories on more than 20 nights since Aug. 28, including CCD B photometry for a few hours each night, with a time resolution of 4-5 min, showing that the star’s average brightness declined at a rate of ∼ 0.01 mag/day. So far, more than a thousand exposures have been analyzed. On every night, the star displayed erratic eruptions on timescales of 1-2 hr, with a typical amplitude of 0.1 mag. This variation is reflected in the power spectrum of the whole data set through a QPO-like feature at frequencies close to 0.5 hr^{−1}. The power spectrum also reveals a power-law rise toward the low-frequency end of the spectrum. We find no evidence for 4.5- or 5-hr stable periodicity in our data.”

PERIODIC COMET SWIFT-TUTTLE (1992t) G. P. Tozzi and G. Calamai, Osservatorio Astronomico Arcetri, Flo-

rence, report the following magnitudes obtained with the Italian National Infrared Telescope TIRGO (28′′ diaphragm), showing a sharp increase in infrared brightness over 24 hr: Nov. 6.7 UT, J = 8.85 ± 0.05, H = 8.45 ± 0.05, K = 8.23 ± 0.05; Nov. 7.7, 8.25 ± 0.10, 7.90 ± 0.10, 7.80 ± 0.10.

PERIODIC COMET BREWINGTON (1992p) Ephemeris extension from orbital elements on MPC 21081:

1992/93 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 14 9^{h}36.^{m}10 +32^{o}36.^{′}1 1.990 2.315 96.^{o}1 25.^{o}1 15.1
24 9 42.02 +32 31.7 1.937 2.384 104.5 23.6 15.5
Dec. 4 9 44.47 +32 39.7 1.886 2.454 113.6 21.6 16.0
14 9 43.30 +32 58.1 1.844 2.524 123.4 19.0 16.4
24 9 38.48 +33 23.0 1.815 2.594 133.8 15.9 16.9
Jan. 3 9 30.32 +33 48.3 1.803 2.665 144.4 12.4 17.3
13 9 19.54 +34 06.5 1.814 2.735 154.5  8.9 17.8

Total visual magnitudes from H. Mikuz, Ljubljana, Slovenia (0.2-m reflector): Sept. 8.13 UT, 11.2; 26.14, 12.3 (CCD); Oct. 9.14, 14.4.

1992 November 10 Daniel W. E. Green